OpenStudy (anonymous):

if the system {y=X^2 +2x+3 y=kx-1 has exactly one solution, then what does k equal

OpenStudy (anonymous):

Equate the coefficients since both are equal x^2+2x+3=kx-1 power of x in 2x is 1 and power x in kx is 1 so k=2

OpenStudy (anonymous):

@ipm1988 but how does that make it equal?

OpenStudy (anonymous):

look at the equations both are equal to y

OpenStudy (anonymous):

y=x^2+2x+3 and y =kx-1 so since both are equal to y therefore x^2+2x+3=kx-1

OpenStudy (anonymous):

@ipm1988 then how would you get from that to k=2

OpenStudy (raden):

so that has exactly one solution then take D=0 (D=discriminan)!

OpenStudy (anonymous):

@kaileyhedgepeth are you familiar with the concept of comparing coefficients of variables of two equations

OpenStudy (anonymous):

your second equation y=kx-2 can also be written as a quadratic equation even though it is a linear equation

OpenStudy (anonymous):

@ipm1988 can you do this problem step by step please?

OpenStudy (raden):

from x^2+2x+3=kx-1 or x^2+2x+3-kx+1=0 x^2+(2-k)x+4=0 just make it D=0, solve for k! do u know what's that discriminan ?

OpenStudy (anonymous):

@RadEn after that step what do you do and no i dont know what that is

OpenStudy (raden):

well, if given a quadratic equation : ax^2+bx+c=0, so its discriminan value is D=b^2-4ac (remember this)! from info above, if given x^2+(2-k)x+4=0 (it means a=1, b=2-k, c=4) so, D=(2-k)^2 - 4(1)(4) D= 4-4k+k^2-16 D=k^2-4k-12 because D=0 therefor, k^2-4k-12=0 or (k-6)(k+2)=0 zeros value of that factors is k-6=0 --->k=6 k+2=0 ---->k=-2 so, k={-2,6}

OpenStudy (anonymous):

@RadEn but there can only be one answer

OpenStudy (raden):

yea, that's right :) it means, if we subtitute k=-2 or k=6 then y=X^2 +2x+3 and y=kx-1, both equation can has only one solution

OpenStudy (raden):

so, the possible values of k are -2 or 6

OpenStudy (raden):

do u understand ?

OpenStudy (anonymous):

yes thanks!

OpenStudy (raden):

yw